# Leveraging Domain Knowledge for Efficient Reward Modelling in RLHF: A Case-Study in E-Commerce Opinion Summarization

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a dominating strategy in steering Language Models (LMs) towards human values/goals. The key to the strategy is employing a reward model ($\varphi$) which can reflect a latent reward model with humans. While this strategy has proven to be effective, the training methodology requires a lot of human preference annotation (usually of the order of tens of thousands) to train $\varphi$. Such large-scale preference annotations can be achievable if the reward model can be ubiquitously used. However, human values/goals are subjective and depend on the nature of the task. This poses a challenge in collecting diverse preferences for downstream applications. To address this, we propose a novel methodology to infuse domain knowledge into $\varphi$, which reduces the size of preference annotation required. We validate our approach in E-Commerce Opinion Summarization, with a significant reduction in dataset size (just $940$ samples) while advancing the state-of-the-art. Our contributions include a novel Reward Modelling technique, a new dataset (PromptOpinSumm) for Opinion Summarization, and a human preference dataset (OpinPref). The proposed methodology opens avenues for efficient RLHF, making it more adaptable to diverse applications with varying human values. We release the artifacts [^1] Code: github.com/efficient-rlhf. PromptOpinSumm: hf.co/prompt-opin-summ. OpinPref: hf.co/opin-pref for usage under MIT License.
Leveraging Domain Knowledge for Efficient Reward Modelling in RLHF: A Case-Study in E-Commerce Opinion Summarization
Swaroop ${}^{\dagger}$${}^{\clubsuit}$, Tejpalsingh ${}^{\dagger}$${}^{\clubsuit}$, Sri ${}^{\dagger}$${}^{\clubsuit}$, Rupasai ${}^{\dagger}$${}^{\clubsuit}$, Harshad ${}^{\diamondsuit}$, Pushpak ${}^{\clubsuit}$, Suman ${}^{\heartsuit}$, Amey ${}^{\heartsuit}$, Sudhanshu Shekhar ${}^{\heartsuit}$, Muthusamy ${}^{\heartsuit}$, Nikesh ${}^{\heartsuit}$$^{\clubsuit}$omputer Science and Engineering, IIT Bombay, India ${}^{\diamondsuit}$ Franklin Templeton ${}^{\heartsuit}$lipkart, India {swaroopnath, tejpalsingh, sriraghava, rupasai, pb}@cse.iitb.ac.in
${}^{\dagger}$ Equal contribution
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2019); Ouyang et al. (2022) has emerged as a dominant paradigm in steering Language Models (LMs) towards human values. In the context of RLHF, human values are represented by a function ($\varphi$). For an output $Y$ ($=y_{1},y_{2},\cdots,y_{n}$) to some input $X$ ($=x_{1},x_{2},\cdots,x_{m}$), $\varphi$ performs the mapping $(X,Y)\rightarrow r$. The reward function $\varphi$ is latent to humans and manifests in human preferences. Preference Modelling techniques, such as Bradley-Terry model Bradley and Terry (1952), Plackett-Luce models Plackett (1975); Luce (2012) are used to learn $\varphi$ from preference data, of the form: $\mathcal{D}=\{(X,Y_{w},Y_{l})\>|\>Y_{w}\succ Y_{l}\}$.
@@
Figure 1: Human Eval: Pairwise win-tie-loss percentage of Inductive-Bias model (our proposed model) vs. ground truth summary and summary from other models, for Amazon benchmark. We see that our proposed technique (domain knowledge infused reward model $\varphi$) helps Inductive-Bias model achieve summaries which are always preferred more than the competitors.
@@
In contemporary works Ziegler et al. (2019); Bai et al. (2022a); Ouyang et al. (2022); Rafailov et al. (2023), the reward functions are parameterized as LMs themselves. Thus the raw text, ($X$, $Y_{w}$) and ($X$, $Y_{l}$) are directly fed to $\varphi$. Such a formulation necessitates large-scale human preference data, to train the large LM (typically millions/billions of parameters). Typically the size of $\mathcal{D}$ varies from $20K$ Ziegler et al. (2019); Nakano et al. (2021); Bai et al. (2022a) to $>200K$ Ethayarajh et al. (2022). Such a large-scale annotation might be achievable if the trained $\varphi$ can be used ubiquitously, irrespective of the nature of downstream application. However, human values are subjective in nature Jiang et al. (2022); Sorensen et al. (2023). For instance, hallucination would be desired in the task of Creative Writing, but undesired in Question-Answering. This means that depending on the context/downstream application, the reward function $\varphi$ can have varying characteristics. Collecting human preferences for all such downstream applications would be impractical.
@@
Motivated to resolve this need, we propose a novel methodology for reward model training. We draw on the insight that $\varphi$ is dependent on the downstream application, and hence, can utilize its domain knowledge. Specifically, $\varphi$ lies in a low-dimensional manifold, whose dimensions can be deduced using domain knowledge. Such an inductive bias reduces the amount of samples [^2] An example: For a function, $f:(x_{1},x_{2},x_{3},\cdots,x_{m})\rightarrow y$, assuming that $f$ is a linear combination of $x_{i}$ (Linear Regression) reduces the training data requirement. Whereas, assuming no functional form (Feed-Forward Neural Network), would lead to more data requirement. needed to learn the parameters of $\varphi$, while increasing bias [^3] Bias of a Machine Learning model is the expected difference between its output and the true value. Say $f^{*}:X\rightarrow y$ is a function, and $\hat{f}$ is an estimate of $f^{*}$, from some training data, then bias is represented as: $\mathbb{E}[\hat{f}(X)-f^{*}(X)]$. of $\varphi$. Concretely our hypothesis is: An inductive bias infused $\varphi$ can help achieve alignment with human values for a task, with modest human preference annotations. We experimentally prove this in the domain of E-Commerce Opinion Summarization Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b) – the task of summarizing user reviews for a product. First, we show that it helps advance the state-of-the-art in Opinion Summarization (Section 5). Second, we analyse how our approach helps the LM achieve alignment with human values for Opinion Summarization (Section 6).
@@
    A novel Reward Modelling technique for RLHF, which leverages Domain Knowledge, to achieve alignment with human values, while significantly reducing human preference annotation. In the domain of Opinion Summarization, we achieve alignment while reducing the dataset size by >21×>21$\times$> 21 × (as compared to the smallest publicly available preference data [^4] The smallest publicly available preference data is not in the domain of Opinion Summarization.). Our approach advances the state of the art: humans prefer our models’ outputs $>68\%$ over the SOTA.
@@
    A new dataset, called PromptOpinSumm, including reviews and summaries for $25763$ products ($229521$ summaries), to train and validate Opinion Summarizer models.
  3. 3.
    A new human preference dataset, called OpinPref, in the domain of Opinion Summarization, consisting of $940$ instances.
@@ 2 Related Works
Steering Language Models (LMs) towards human goals: Steering LMs towards human goals/values refers to the task of training LMs to generate text which is more aligned with human values, such as ‘text should not have harmful content’, ‘it should be polite’, etc. Such a task necessitates a human presence in the training loop of the these LMs. In recent times, Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2019); Askell et al. (2021); Bai et al. (2022a); Ouyang et al. (2022); Liu et al. (2022) has emerged as a viable solution – by incorporating Reward Models, which reflect latent reward models wthin humans, into the training pipeline. These reward models are trained on human preference datasets Ziegler et al. (2019); Nakano et al. (2021); Ethayarajh et al. (2022), which are typically of the order of tens of thousands, in size. Dependence on high-quality large-sized preference data is an obstacle for RLHF.
@@
Recently, Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022b); Kim et al. (2023); Lee et al. (2023) has emerged as an alternative. It attempts to reduce the dependence on human preference datasets, by using Large LMs as preference data generators. While this is a scalable approach to steering LMs, there is no guarantee that the preference dataset generated by Large LMs actually reflects human goals – it is still a very open question. In our work, we propose a different solution – which promises to use human preference data, but provides a way to drastically reduce the required size. To the best of our knowledge, we are the first to attempt this.
@@
Opinion Summarization: Opinion Summarization Hu and Liu (2004); Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b) is the task of summarizing user reviews. Specifically, we look at E-Commerce Opinion Summarization, where user reviews are on products. These reviews contain aspects of the product, and users’ sentiments/opinions towards those aspects. Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in E-Commerce Opinion Summarization have used Self-Supervised training methodology. In this context, self-supervision refers to picking one of the $N$ available reviews as a summary, commonly called pseudo-summary, and training the model on the remaining $N-1$ reviews to generate the pseudo-summary. The theme of solutions Chu and Liu (2018); Bražinskas et al. (2020); Siledar et al. (2023b, a) have mostly centered around Supervised Learning. The core problem has always been getting good synthetic datasets for training. More recently, Prompting Bhaskar et al. (2023) has been explored to solve the task. Bhaskar et al. (2023) move away from making a better synthetic dataset generation pipeline, and test Gpt-$3.5$ for Opinion Summarization.
@@
We do not propose a new synthetic dataset generation methodology. Rather, we generate training data using an open-source Large LM (Mistral-$7$B), to test our hypothesis. To the best of our knowledge, we are the first to propose such a dataset for training Opinion Summarizers. Such an approach has been explored in earlier works Wang et al. (2023); Taori et al. (2023); Peng et al. (2023) for general domain text generation. Taori et al. (2023) fine-tune LLaMA-$7$B Touvron et al. (2023a) using Instruction-Tuning dataset generated using Gpt-$3$. Peng et al. (2023) fine-tune LLaMA-$7$B using a dataset generated by Gpt-$4$.
@@ 3 Dataset
Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in Opinion Summarization have used Self-Supervised training methodology. In the context of Opinion Summarization, self-supervision refers to picking one of the $N$ available reviews as a summary, commonly called pseudo-summary, and training the model on the remaining $N-1$ reviews to generate the pseudo-summary. Among early works in this direction, Bražinskas et al. (2020) randomly selected a review as a pseudo-summary to construct a training sample. While, more recently, Siledar et al. (2023a) propose a sophisticated synthetic data creation pipeline for self-supervised opinion summarization.
Although these self-supervision datasets have helped further State-of-the-Art in Opinion Summarization, the trained models have several pitfalls. A few shortcomings are: (a) the generated opinion summaries are in first-person, (b) the review chosen as pseudo-summary might not cover all aspects and opinions, etc. To overcome these shortcomings, we move away from the self-supervised training regime in our work, and propose a new dataset. In the rest of this Section, we describe (a) PromptOpinSumm – a new dataset to train Opinion Summarization models, (b) the benchmarks we used to test our proposed technique, and (c) OpinPref – human preference annotated dataset for Opinion Summarization.
@@ 3.1 PromptOpinSumm Dataset
We prompt the Mistral-$7$B model Jiang et al. (2023) to generate opinion summary given reviews of a product. We also tried other open-source LLMs available at the time of the work, such as LLaMA2222-$7$B, LLaMA2222-$13$B Touvron et al. (2023b), Vicuna-$7$B, Vicuna-$13$B Chiang et al. (2023), Zephyr-$7$B Tunstall et al. (2023). However, we found that Mistral-$7$B leads to better summaries. We limit ourselves to open-source models mainly to limit the expenditure. Appendix E includes an examples along with qualitative analysis. We use the Amazon dataset He and McAuley (2016), which has reviews for $\sim 180k$ products. We sample reviews for $20763$ products for train set and $5000$ products for validation set to prompt the Mistral-$7$B model. Specifically, we prompt the model to generate opinion summaries of $3$ different qualities – Good (codenamed Good-Sum in the rest of the paper), Slightly Bad (codenamed SBad-Sum in the rest of the paper) and Very Bad (codenamed VBad-Sum in the rest of the paper). Again, per quality, we generate multiple opinion summaries ($3$ at most). We provide a reasoning of generating multiple summaries of different qualities in the discussion of our technique (Section 4.2). We generate a total of $184620$ summaries for train set and $44901$ summaries for validation set. Table 1 includes summary statistics of the generated dataset.
Split Characteristic $\mu$ $\sigma$
train # reviews $13.24$ $10.07$
# summaries $8.90$ $0.34$
# words in review $49.0$ $10.78$
# words in summary $78.28$ $34.45$
validation # reviews $10.53$ $6.80$
# summaries $8.98$ $0.16$
# words in review $48.65$ $10.63$
# words in summary $74.26$ $34.27$
Table 1: Statistics of PromptOpinSumm dataset.
3.2 Datasets for Evaluation
We test all our models on $3$ popular benchmarks: Amazon Bražinskas et al. (2020), Oposum+ Amplayo et al. (2021), and Flipkart Siledar et al. (2023b). Amazon dataset has reviews for $32$ products from $4$ domains: “electronics”, “home & kitchen”, “personal care”, and “clothing, shoes & jewellery”. Oposum+ dataset has reviews for $60$ products from $6$ domains: “laptop bags”, “bluetooth headsets”, “boots”, “keyboards”, “television”, and “vacuums”. Flipkart dataset has reviews for $147$ products from $3$ domains: “laptops”, “mobiles”, and “tablets”. Table 2 includes relevant statistics of these benchmarks.
Characteristic Oposum+ Amazon Flipkart
# domains $6$ $4$ $3$
# products $60$ $32$ 147
# reviews per product $10$ $8$ $10$
# summaries per product $3$ $3$ $1$
Table 2: Statistics of the benchmark datasets.
While we acknowledge the widespread use of these benchmarks, we notice several shortcomings. We highlight them in Appendix B. Due to these shortcomings, in our analysis, we do not rely much on overlap-based evaluations, such as Rouge. Rather we rely on human evaluations.

### v2

# Leveraging Domain Knowledge for Efficient Reward Modeling in RLHF: A Case-Study in E-Commerce Opinion Summarization

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a dominating strategy in aligning Language Models (LMs) with human values/goals. The key to the strategy is learning a reward model ($\varphi$), which can reflect the latent reward model of humans. While this strategy has proven effective, the training methodology requires a lot of human preference annotation (usually in the order of tens of thousands) to train $\varphi$. Such a large-scale annotation is justifiable when it’s a one-time effort, and the reward model is universally applicable. However, human goals are subjective and depend on the task, requiring task-specific preference annotations, which can be impractical to fulfill. To address this challenge, we propose a novel approach to infuse domain knowledge into $\varphi$, which reduces the amount of preference annotation required (21×21$\times$21 ×), omits Alignment Tax, and provides some interpretability. We validate our approach in E-Commerce Opinion Summarization, with a significant reduction in dataset size (to just $940$ samples) while advancing the SOTA ($\sim 4$ point Rouge-$L$ improvement, $68\%$ of times preferred by humans over SOTA). Our contributions include a novel Reward Modeling technique and two new datasets: PromptOpinSumm (supervised data for Opinion Summarization) and OpinPref (a gold-standard human preference dataset). The proposed methodology opens up avenues for efficient RLHF, making it more adaptable to applications with varying human values. We release the artifacts [^1] Code: github.com/efficient-rlhf. PromptOpinSumm: hf.co/prompt-opin-summ. OpinPref: hf.co/opin-pref for usage under MIT License.
Leveraging Domain Knowledge for Efficient Reward Modeling in RLHF: A Case-Study in E-Commerce Opinion Summarization
Swaroop Nath, Tejpalsingh Siledar, Sri Raghava, Rupasai Rangaraju, Harshad Khadilkar, Pushpak Bhattacharyya, Suman Banerjee, Amey Patil, Sudhanshu Shekhar Singh, Muthusamy Chelliah, Nikesh Garera Computer Science and Engineering, IIT Bombay, India Flipkart, India swaroopnath@cse.iitb.ac.in
Equal contribution
@@ 1 Introduction
[…]
@@
Figure 1: Human Eval: Pairwise win-tie-loss percentage of Inductive-Bias model (our proposed model) vs. ground truth summary and summary from other models, for Amazon benchmark. We see that our proposed approach (infusing domain knowledge into $\varphi$ to reap benefits of RLHF with modest human preference data) helps Inductive-Bias model achieve summaries which are always preferred (Section 5.2).
@@
In contemporary works Ziegler et al. (2019); Bai et al. (2022a); Ouyang et al. (2022); Rafailov et al. (2023), the reward functions are Large LMs (LLMs; pretrained Transformers) themselves. The text data, ($X$, $Y_{w}$) and ($X$, $Y_{l}$) are directly fed to $\varphi$, for training. Such a formulation necessitates large-scale human preference data to train the LLM (millions/billions of parameters). Typically the size of $\mathcal{D}$ varies from $20K$ Nakano et al. (2021); Bai et al. (2022a) to $>200K$ Ethayarajh et al. (2022). Such a large-scale annotation is justifiable when it’s a one-time effort, and the trained $\varphi$ is universally applicable, irrespective of the nature of the downstream task. However, human values are subjective Jiang et al. (2022); Sorensen et al. (2023). For instance, hallucination would be desired in Creative Writing, but not in Question-Answering. This means that depending on the downstream task, the reward function $\mathbf{\varphi}$ must have varying characteristics. Collecting human preferences for all such tasks is impractical.
@@
Motivated to resolve this need, we propose a novel reward modeling methodology, significantly reducing preference data requirements. We draw on the insight that $\varphi$ is dependent on the downstream task and, hence, can utilize its task/domain [^3] We use task and domain interchangeably in the paper. knowledge. Specifically, $\varphi$ lies in a low-dimensional manifold, whose dimensions can be deduced using domain knowledge. Such an inductive bias reduces the number of samples [^4] An example: For a function, $f:(x_{1},x_{2},x_{3},\cdots,x_{m})\rightarrow y$, assuming that $f$ is a linear combination of $x_{i}$ (Linear Regression) reduces the training data requirement. Assuming no functional form (Feed-Forward Neural Network) would require more data. needed to train $\varphi$. Concretely, our hypothesis is: An inductive bias infused $\varphi$ can help achieve alignment with human values for a task, with modest human preference annotations. Specifically, we say that $\varphi_{\tau}$ (reward model for a domain $\tau$) can be modelled by some numeric features $v_{1},v_{2},\cdots,v_{n}$. These $n$ features fully characterize [^5] Example of such characterization: Features like fluency, coherence, etc. can characterize text generated by an LLM. the outputs from the LLM on some input. Thus, instead of training $\varphi_{\tau}$ on the text data ($\{(X,Y_{w},Y_{l})\>|\>Y_{w}\succ Y_{l}\}$), we use the $n$ features. Such a formulation for $\varphi$ brings interpretability—which features influence human preference the most (Section 6), and is free from Alignment Tax (degradation of language capabilities of an LLM post reward modeling; Bai et al. (2022a)) as we do not use an LLM to model $\varphi$.
We experimentally prove our hypothesis in the domain of E-Commerce Opinion Summarization Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b)—the task of summarizing user reviews for a product. In addition to advancing SOTA, we also analyze how our approach helps the model achieve alignment with human values for Opinion Summarization (Section 6).
@@
    A novel Reward Modeling technique for RLHF, which leverages Domain Knowledge to achieve alignment with human values while significantly reducing human preference annotation. In the domain of Opinion Summarization, we achieve alignment while reducing [^6] As compared to the smallest publicly available preference data. The smallest publicly available preference data is not in the domain of Opinion Summarization. the dataset size by $>\mathbf{21\times}$. Our approach advances SOTA: at least $\sim\mathbf{4}$-point Rouge-$L$ improvement (Tables 1, 4 and 5; Section 5.2), and humans prefer our models’ outputs $\mathbf{68}\%$ over SOTA (Figure 1; Section 5.2).
@@
    Two new datasets: PromptOpinSumm and OpinPref. PromptOpinSumm includes reviews and summaries for $25763$ products ($229521$ summaries), for training and validation. OpinPref is a gold-standard human preference dataset (with $940$ instances) in the domain of Opinion Summarization.
@@ 2 Related Works
Steering Language Models (LMs) towards human goals: Steering LMs towards human goals/values refers to the task of training LMs to generate text which is more aligned with human values, such as ‘text should not have harmful content’, ‘it should be polite’, etc. Such a task necessitates a human presence in the training loop of these LMs. In recent times, Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2019); Askell et al. (2021); Bai et al. (2022a); Ouyang et al. (2022); Liu et al. (2022) has emerged as an effective solution—by incorporating Reward Models, which reflect latent reward models within humans, into the training pipeline. These reward models are trained on human preference datasets Ziegler et al. (2019); Nakano et al. (2021); Ethayarajh et al. (2022), which are typically of the order of tens of thousands, in size. Dependence on high-quality, large-sized preference data is an obstacle for RLHF.
@@
Recently, Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022b); Kim et al. (2023); Lee et al. (2023) has emerged as an alternative. It attempts to reduce the dependence on human preference datasets by using Large LMs (LLMs) as preference data generators. While this is a scalable approach to steering LMs, there is no guarantee that the preference dataset generated by LLMs reflects human goals. In our work, we propose a different solution, which promises to use human preference data but provides a way to reduce the required size drastically. To the best of our knowledge, we are the first to attempt this.
@@
Opinion Summarization: Opinion Summarization Hu and Liu (2004); Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b) is the task of summarizing user reviews. Specifically, we look at E-Commerce Opinion Summarization, where user reviews are on products. These reviews contain aspects of the product and users’ sentiments/opinions towards those aspects. Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in E-Commerce Opinion Summarization have used Self-Supervised training methodology. In this context, self-supervision refers to picking one of the $N$ available reviews as a summary, commonly called pseudo-summary, and training the model on the remaining $N-1$ reviews to generate the pseudo-summary. The theme of solutions Chu and Liu (2018); Bražinskas et al. (2020); Siledar et al. (2023b, a) have mostly centered around Supervised Learning. The core problem has always been getting good synthetic datasets for training. More recently, Prompting Bhaskar et al. (2023) has been explored to solve the task. Bhaskar et al. (2023) move away from making a better synthetic dataset generation pipeline and test Gpt-$3.5$ for Opinion Summarization.
@@
We do not propose a new synthetic dataset generation methodology. Rather, we generate training data using an open-source LLM (Mistral-$7$B), to test our hypothesis. To the best of our knowledge, we are the first to propose such a dataset for training Opinion Summarizers. Such an approach has been explored for Generic Text Summarization Wang et al. (2023); Taori et al. (2023); Peng et al. (2023). Taori et al. (2023) fine-tune LLaMA-$7$B Touvron et al. (2023a) using Instruction-Tuning dataset generated using Gpt-$3$. Peng et al. (2023) fine-tune LLaMA-$7$B using a dataset generated by Gpt-$4$.
@@ 3 Dataset
Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in Opinion Summarization have used Self-Supervised training methodology, where $N-1$ reviews are used as input, and the left out review is used as a pseudo-summary (Section 2). Although these self-supervision datasets have helped further Opinion Summarization research, the approach has several shortcomings: the summaries always present a one-person rather than the consensus view, the summaries are reviews and might not have good coverage of aspects and opinions, etc. We move away from self-supervision to overcome these shortcomings and propose a new dataset. In the rest of this Section, we describe (a) PromptOpinSumm: a new dataset to train Opinion Summarizers, (b) the benchmarks we used for evaluation, and (c) OpinPref: gold-standard preference dataset for Opinion Summarization.
@@ 3.1 PromptOpinSumm Dataset
We prompt the instruction-tuned Mistral-$7$B model Jiang et al. (2023) to generate an opinion summary given product reviews. We also tried other open-source LLMs available at the time of the work, such as LLaMA2222-$7$B, LLaMA2222-$13$B Touvron et al. (2023b), Vicuna-$7$B, Vicuna-$13$B Chiang et al. (2023), Zephyr-$7$B Tunstall et al. (2023). However, we found that Mistral-$7$B leads to better summaries. We limit ourselves to open-source models due to cost. Appendix G includes examples and qualitative analysis. We use the Amazon dataset He and McAuley (2016), which has reviews for $\sim 180k$ products. We randomly sample reviews for $20763$ products for train set and $5000$ products for validation set. Specifically, we prompt the model to generate opinion summaries of $3$ different qualities: Good (codenamed Good-Sum), Slightly Bad (codenamed SBad-Sum), and Very Bad (codenamed VBad-Sum). We generate multiple opinion summaries ($3$ at most) per quality. We provide reasoning for generating multiple summaries of different qualities in the extended discussion of our approach (Appendix B). We generate $184620$ summaries for train set and $44901$ summaries for validation set (see Appendix G).
3.2 Benchmarks for Evaluation
We use $9$ Opinion Summarization benchmarks for evaluation. $3$ of these benchmarks are the Amazon test set (Bražinskas et al. (2020), codenamed Amazon), the Oposum+ test set (Amplayo et al. (2021), codenamed Oposum+) and the Flipkart test set (Siledar et al. (2023b), codenamed Flipkart). Amazon has reviews for $32$ products from $4$ domains, Oposum+ has reviews for $60$ products from $6$ domains and Flipkart has reviews for $147$ products from $3$ domains.
Although these $3$ benchmarks have been used widely, they have several shortcomings. For instance, Amazon was developed by asking annotators to write a summary in first-person point-of-view. This causes problems such as summaries seeming personal rather than consensus opinions (which can include mixed sentiment), incomplete coverage of aspects and opinions, etc. Thus, using such pseudo-summaries for reference-based evaluations (Rouge, BertScore) on such a benchmark is not a correct portrayal of the models’ performances. We highlight the shortcomings in detail in Appendix D. Siledar et al. (2024) recently provided $6$ new benchmarks (Amazon-R, Amazon-Rdq, Oposum-R, Oposum-Rdq, Flipkart-R, Flipkart-Rdq) which are revamped versions (by getting rid of the shortcomings) of the aforementioned $3$ benchmarks. We primarily rely on these $6$ for our conclusions. Appendix D includes more details on domains and summary statistics.